OpenStudy (anonymous):

The Ruiz family is going to build a raised garden bed for planting flowers in. The bed will be 3/4 meters wide and 3 1/2 meters long. What will the area be when the bed is finished? I need the expression

OpenStudy (anonymous):

First, I would change the fractions to decimals. 3/4=.75 1/2=.5 So now we have .75 and 3.5. I assume the garden bed is a rectangle because we are only given two sides. To find the area of a rectangle, take one side and multiply it by the other side. 0.75*3.5=?

OpenStudy (anonymous):

Thank you very much

OpenStudy (anonymous):

Your welcome.

OpenStudy (anonymous):

Also, if the answer choices are in fractions, 0.625=5/8

OpenStudy (anonymous):

Thank you again, I was trying to figure that out